OpenStudy (anonymous):

Passage: Kevin’s family has a history of diabetes. The probability that Kevin would inherit this disease is 0.75. Kevin decides to take a test to check if he has the disease. The accuracy of this test is 0.85.We now know that the test predicted that Kevin does not have the disease. Using this information, calculate the probability that Kevin does not have diabetes after the test was taken.

OpenStudy (anonymous):

help

OpenStudy (kropot72):

 P(false negative) = 0.75 * 0.15 = 0.1125 P(true negative) = 0.25 * 0.85 = 0.2125 P(negative result) = 0.1125 + 0.2125 = 0.325 P(not diabetic after test) = 0.2125/0.325 = 0.654